# Anchors Aweigh! Sail for Optimal Unified Multi-Modal Representations

### v2

@@ 4.2 Experiments with real-world dataset
[…]
Downstream tasks.
We perform evaluations in zero-shot binary and multi-class classification tasks, One-to-One cross-modal retrieval, and Two-to-One cross-modal retrieval. For classification tasks, we use a Multi-Layer Perceptron (MLP) to perform sarcasm detection as a binary classification and speaker classification with $23$ multi-class categories. In particular, MLP is trained on embeddings in a single modality (denoted by Tr in Table 2) and accuracy is evaluated on another modality (denoted by Ev in Table 2). In retrieval tasks, we measure the accuracy of correct retrieval. For the One-to-One case, we retrieve data sample in different modality by choosing the closest embedding from a single input embedding, while for the Two-to-One case we choose the closest embedding from the centroid of two input embeddings in two modalities. We denote input and target modalities with $\rightarrow$ in Table 1.
@@
Table 2: Accuracy results for Sarcasms and Speakers. ($\mathcal{V}$: video, $\mathcal{A}$: audio, $\mathcal{T}$: text). Asterisks denote accuracy evaluated in different settings.
[…]
@@
Table 2 presents results for sarcasm detection and speaker classification tasks, where Sar-1 indicates Top-1 accuracy for sarcasm, and Spk-$k$, $k=1,3,5$ represent Top-k accuracies for speaker classification. It is important to highlight that CentroBind and FABind are trained on a single modality (Tr) and evaluated on a different modality (Ev) in a zero-shot setting, which can effectively measure the ability of multi-modal alignment. In this experiment, CentroBind consistently outperforms FABind and UniBind across all pairs of train and evaluation modalities, which can be distributed to CentroBind generally learning a better unified embedding space than FABind. UniBind performs poorly in the zero-shot cross-modal experiment, which we believe is due to its insufficient multi-modal alignment. Since UniBind utilizes LLM-augmented descriptions for each modality and binds other encoders to these descriptions, multi-modal alignment may fail if the descriptions are dispersed across the embedding space. As analyzed in Section 2.3 and Section 3.2, these results highlight the CentroBind’s ability to preserve intra and shared information among modalities, which are useful in unknown downstream tasks. Moreover, the zero-shot setting verifies the multi-modal alignment of CentroBind.
@@ 5 Conclusions
In this paper, we analyze the limitations of fixed-anchor-binding (FABind) methods, including their over-reliance on a single anchor modality and their inability to capture both intra-modal and shared information among non-anchored modalities. To address these shortcomings, we propose adaptive anchor binding methods such as CentroBind, which align multi-modal embeddings to centroid-based anchors, removing the need for a fixed anchor modality. Our approach generalizes and extends methods like ImageBind. We also provide a theoretical analysis showing that CentroBind effectively captures both intra-modal and shared inter-modal information. Experiments on synthetic and real-world datasets demonstrate that CentroBind significantly outperforms FABind, offering a robust unified representation space and validating our theoretical insights.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Multi-modal learning
Typically, these models integrate uni-modal features extracted by modality-specific encoders (Seichter et al., 2021; Nagrani et al., 2021; Wu et al., 2022; Wang et al., 2020; Peng et al., 2022). For instance, Madaan et al. (2024) introduce inter- and intra-modality modeling frameworks that treat the target as a composition of multiple modalities. Similarly, Du et al. (2023) propose a late-fusion approach for supervised multi-modal tasks, demonstrating that insufficient feature extraction from individual modalities negatively affects the model’s generalization ability. Additionally, Zhang et al. (2024) address joint optimization by alternating between uni-modal learning scenarios and integrating modality-specific encoders with a unified head shared across all modalities.
Multi-view learning is closely related to multi-modal learning. Early work by Tian et al. (2020a) study multi-view representation learning through predictive and contrastive approaches across multiple modalities, demonstrating the efficacy of contrastive learning in multi-view settings. Building on these insights, numerous subsequent studies have integrated contrastive learning into multi-view tasks. For example, Xing et al. (2019) propose an approach that adaptively combines image-text representations for few-shot learning; however, this method still relies on labeled data. Some research leverages anchors or a fused modality as the anchor representation. Zeng et al. (2021) introduce a unified prototype representation to address cross-modal retrieval imbalance by employing a reconstruction loss, but its reliance on unified prototypes as anchors restricts the capture of inter-view information across modalities. Meanwhile, Huang et al. (2023) develop a tri-modal alignment pre-training task—extending text-video alignment to include a fused modality using pairwise contrastive learning; however, it does not explicitly handle intra-view learning. Furthermore, Wang et al. (2022b) present a cross-modal data augmentation technique for image-text multi-view learning, randomly replacing visually grounded words with diverse image patches to increase data variety and encourage token-level interaction across modalities. Nevertheless, this approach still overlooks intra-view information, indicating the need for methods that jointly consider both inter- and intra-view representations.
@@ A.2 Multi-modal alignment
Multi-modal learning addresses four key challenges (Liang et al., 2024c; Baltrušaitis et al., 2018; Liang et al., 2024d): managing interactions among redundant, unique, and synergistic features (Dumas et al., 2017; Liang et al., 2024a, b), aligning fine-grained and coarse-grained information (Wang et al., 2023, 2024a), reasoning across diverse features (Yang et al., 2023), and integrating external knowledge (Shen et al., 2022; Lyu et al., 2024). Among these challenges, multi-modal alignment is one of the core challenges that many researchers aim to solve.
A common method in multi-modal alignment is using cross-modal alignment by using attention mechanisms between pairwise modalities, such as vision-language (Tan & Bansal, 2019) and vision-language-audio (Tsai et al., 2019). Another effective approach is leveraging graph neural networks to align multi-modal datasets (Yang et al., 2021; Wilf et al., 2023). For instance, Yang et al. (2021) transform unaligned multi-modal sequence data into nodes, with edges capturing interactions across modalities over time. Wilf et al. (2023) build graph structures for each modality—visual, textual, and acoustic—and create edges to represent their interactions.
To enhance the generalizability of cross-modal representations, Xia et al. (2024) employ a unified codebook approach, facilitating a joint embedding space for visual and audio modalities. Another prominent method achieves cross-modal alignment by leveraging large collections of image-text pairs, making it a widely adopted strategy in multi-modal learning (Radford et al., 2021; Zhang et al., 2022; Guzhov et al., 2022; Zhou et al., 2023).
A.3 Binding methods
@@ Appendix B Proofs
=$\overset{\rm(b)}{=}\frac{M}{\tau}{\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,k})-|\mathcal{I}_{B}|\log\sum_{j\in\mathcal{I}_{B}}\exp\left(\frac{M{\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,j})}{\tau|\mathcal{I}_{B}|}\right)-|\mathcal{I}_{B}|\log C_{\mathcal{F},k,i}$
@@
$I_{\rm NCE}\left({\bf A};f_{i}({\bf X}_{i})\;\middle|\;\tilde{\tau}\right)$ $\geq\frac{1}{|\mathcal{I}_{B}|}\sum_{l=1}^{M}I_{\rm NCE}\left(f_{l}({\bf X}_{l}^{\prime});f_{i}({\bf X}_{i})\;\middle|\;\frac{\tilde{\tau}M}{|\mathcal{I}_{B}|}\right)-\frac{1}{|\mathcal{I}_{B}|}\sum_{k=1}^{|\mathcal{I}_{B}|}\log C_{\mathcal{F},k,i},$ (26)
@@
Appendix C Experiment
@@ C.1 Experiments with synthetic datasets
$p_{\bf Z}({\boldsymbol{z}})$ $=\prod_{y=1}^{K}\pi_{y}\mathcal{N}({\boldsymbol{z}};\hbox{\boldmath$\mu$}_{y},\hbox{\boldmath$\Sigma$}_{y}),$ (27)
[…]
@@
Figure 5: Accuracy as a measure of the representation space quality. Abbreviation: ${\bf X}_{i}$-B or CB: applying FABind with anchor ${\bf X}_{i}$ or applying CentroBind; acc$({\bf Z}_{i})$ or acc(All): accuracy of ${\bf Z}_{i}$ or of concatenated embeddings $({\bf Z}_{1},\cdots,{\bf Z}_{M})$; (rnd): if random backbones are used.
@@
The plot also reveals a crossover point where the loss curves intersect. We believe this occurs due to the number of InfoNCE losses optimized by CentroBind and FABind. Specifically, with $M$ modalities, CentroBind minimizes $M$ InfoNCE losses, while FABind minimizes $M-1$ InfoNCE losses. This results in a smaller loss for FABind when the encoders are well-trained, which explains the crossover point observed in Figure 7.
@@
[…]
@@
Figure 8: Experiment results with synthetic dataset of $M=6$ modalities. Abbreviation: ${\bf X}_{i}$-B or CB: applying FABind method to backbones with anchor ${\bf X}_{i}$ or applying CentroBind; acc$({\bf Z}_{i})$ or acc(All): accuracy of ${\bf Z}_{i}$ or of concatenated embeddings $({\bf Z}_{1},\cdots,{\bf Z}_{M})$; (rnd): if random backbones are used for ${\bf X}_{i}$-B or CB.
@@
Figure 9: Experiment results with synthetic dataset of $M=8$ modalities. Abbreviation: ${\bf X}_{i}$-B or CB: applying FABind method to backbones with anchor ${\bf X}_{i}$ or applying CentroBind; acc$({\bf Z}_{i})$ or acc(All): accuracy of ${\bf Z}_{i}$ or of concatenated embeddings $({\bf Z}_{1},\cdots,{\bf Z}_{M})$; (rnd): if random backbones are used for ${\bf X}_{i}$-B or CB.
@@
Figure 10: Comparison of other dynamic anchor generation methods. (a) and (b): Modal qualities are set to $(0.2,0.2,0.2,1)$. (c) and (d): Modal qualities are set to $(0.2,0.2,0.8,0.8)$. Abbreviation: ${\bf X}_{i}$-B or CB: applying FABind method to backbones with anchor ${\bf X}_{i}$ or applying CentroBind; WAB: weighted average for dynamic anchor with weight identical to the predefined quality for each modality; RB+Intra: randomly choosing a modality for a dynamic anchor in every iteration and intra information learning; RB: randomly choosing a modality for a dynamic anchor in every iteration; MB: coordinate-wise median for dynamic anchors; acc$({\bf Z}_{i})$ or acc(All): accuracy of ${\bf Z}_{i}$ or of concatenated embeddings $({\bf Z}_{1},\cdots,{\bf Z}_{M})$; (ran): if random backbones are used.
C.2 Experiments with real-world datasets

### v3

@@ 4.2 Experiments with real-world datasets
One-to-One Two-to-One Method Retrieval Top-1 Top-5 Top-10 Method Retrieval Top-1 Top-5 Top-10 FABind $\mathcal{V}$ $\rightarrow$ $\mathcal{T}$ 0.446 0.719 0.822 FABind $\mathcal{V}$, $\mathcal{A}$ $\rightarrow$ $\mathcal{T}$ 0.309 0.665 0.781 CentroBind 0.483 0.764 0.850 FABind $\mathcal{A}$ $\rightarrow$ $\mathcal{T}$ 0.077 0.238 0.367 CentroBind 0.745 0.957 0.978 CentroBind 0.233 0.517 0.678 FABind $\mathcal{T}$ $\rightarrow$ $\mathcal{V}$ 0.812 0.946 0.978 FABind $\mathcal{T}$, $\mathcal{A}$ $\rightarrow$ $\mathcal{V}$ 0.180 0.401 0.513 CentroBind 0.591 0.839 0.909 FABind $\mathcal{A}$ $\rightarrow$ $\mathcal{V}$ 0.058 0.154 0.226 CentroBind 0.388 0.646 0.768 CentroBind 0.052 0.184 0.284 FABind $\mathcal{T}$ $\rightarrow$ $\mathcal{A}$ 0.201 0.438 0.584 FABind $\mathcal{T}$, $\mathcal{V}$ $\rightarrow$ $\mathcal{A}$ 0.099 0.257 0.364 CentroBind 0.290 0.572 0.706 FABind $\mathcal{V}$ $\rightarrow$ $\mathcal{A}$ 0.051 0.155 0.223 CentroBind 0.232 0.490 0.625 CentroBind 0.054 0.175 0.258
Downstream tasks with MUStARD.
We perform evaluations in zero-shot binary and multi-class classification tasks, One-to-One, and Two-to-One cross-modal retrieval. For classification tasks, we use a Multi-Layer Perceptron (MLP) to perform sarcasm detection as a binary classification and speaker classification with $23$ multi-class categories. In particular, MLP is trained on embeddings in a single modality (denoted by Tr in Table 2) and accuracy is evaluated on another modality (denoted by Ev in Table 2). In retrieval tasks, we measure the accuracy of correct retrieval. For the One-to-One case, we retrieve data sample in different modality by choosing the closest embedding from a single input embedding, while for the Two-to-One case we choose the closest embedding from the centroid of two input embeddings in two modalities. We denote input and target modalities with $\rightarrow$ in Table 1.
@@
Table 2 presents results for sarcasm detection and speaker classification tasks, where Sar-1 indicates Top-1 accuracy for sarcasm, and Spk-$k$, $k=1,3,5$ represent Top-k accuracies for speaker classification. It is important to highlight that CentroBind and FABind are trained on a single modality (Tr) and evaluated on a different modality (Ev) in a zero-shot setting, which can effectively measure the ability of multi-modal alignment.
Table 2: Zero-shot accuracy results. ($\mathcal{V}$: video, $\mathcal{A}$: audio, $\mathcal{T}$: text). Asterisks: accuracy evaluated in different settings.
Method Tr, (Ev) Sar-1 Spk-1 Spk-3 Spk-5 FABind $\mathcal{V}$, ($\mathcal{T}$) 0.706 0.378 0.614 0.730 UniBind 0.544 0.170 0.328 0.478 AudioCLIP 0.501 0.096 0.258 0.388 ViT-Lens 0.506 0.097 0.343 0.449 CentroBind 0.716 0.474 0.736 0.836 FABind $\mathcal{A}$, ($\mathcal{T}$) 0.648 0.186 0.455 0.577 UniBind 0.628 0.220 0.399 0.501 AudioCLIP 0.486 0.094 0.214 0.322 ViT-Lens 0.484 0.077 0.214 0.313 CentroBind 0.691 0.290 0.546 0.714 FABind $\mathcal{T}$, ($\mathcal{V}$) 0.572 0.243 0.445 0.630 UniBind 0.484 0.129 0.262 0.404 AudioCLIP 0.506 0.158 0.345 0.461 ViT-Lens 0.502 0.168 0.323 0.423 CentroBind 0.694 0.368 0.670 0.791 FABind $\mathcal{A}$, ($\mathcal{V}$) 0.623 0.228 0.484 0.628 UniBind 0.567 0.199 0.367 0.514 AudioCLIP 0.503 0.209 0.384 0.496 ViT-Lens 0.500 0.149 0.332 0.451 CentroBind 0.683 0.243 0.475 0.632 FABind $\mathcal{V}$, ($\mathcal{A}$) 0.604 0.255 0.472 0.636 UniBind 0.506 0.126 0.280 0.429 AudioCLIP 0.501 0.080 0.199 0.326 ViT-Lens 0.533 0.219 0.438 0.575 CentroBind 0.626 0.326 0.548 0.703 FABind $\mathcal{T}$, ($\mathcal{A}$) 0.534 0.241 0.509 0.635 UniBind 0.514 0.091 0.248 0.365 AudioCLIP 0.477 0.088 0.309 0.439 ViT-Lens 0.475 0.070 0.214 0.329 CentroBind 0.655 0.346 0.610 0.741
In this experiment, CentroBind consistently outperforms FABind and UniBind across all pairs of train and evaluation modalities, which can be distributed to CentroBind generally learning a better unified embedding space than FABind. UniBind performs poorly in the zero-shot cross-modal experiment, which we believe is due to its insufficient multi-modal alignment. Since UniBind utilizes LLM-augmented descriptions for each modality and binds other encoders to these descriptions, multi-modal alignment may fail if the descriptions are dispersed across the embedding space. As analyzed in Section 2.3 and Section 3.2, these results highlight the CentroBind’s ability to preserve intra and shared information among modalities, which are useful in unknown downstream tasks. Moreover, the zero-shot setting verifies the multi-modal alignment of CentroBind.
Comparison with ImageBind backbone.
We compare the performance of FABind and CentroBind using the ImageBind backbone (Girdhar et al., 2023) to assess whether CentroBind yields additional gains over ImageBind, which was pretrained on large-scale datasets and is optimized primarily for the image modality. As shown in Table 3, CentroBind outperforms ImageBind on the DreamBooth (Ruiz et al., 2023), AVE (Tian et al., 2018), AudioSet (Gemmeke et al., 2017), and UR-FUNNY (Hasan et al., 2019) datasets—even with 1) a strong pretrained backbone, 2) an image-anchored modality, and 3) a bimodal setup. We expect this gap to widen with a weaker backbone, without the image modality, or as additional modalities are introduced, as also evidenced by the synthetic and MUStARD results. These findings align with our analysis of dynamic anchor binding and highlight its effectiveness across nearly all evaluated scenarios. See Appendix C.3 for more discussion.
Table 3: Cross-modal retrieval accuracy. Performance dynamics are shown in Fig 11.
ImageBind CentroBind
Dataset Modalities Top1 Top5 Top1 Top5
DreamBooth $\mathcal{V},\mathcal{T}$ 0.672 0.984 0.719 ($\uparrow 0.047$) 1.000 ($\uparrow 0.016$)
AVE $\mathcal{V},\mathcal{A}$ 0.313 0.649 0.327 ($\uparrow 0.014$) 0.676 ($\uparrow 0.027$)
AudioSet $\mathcal{V},\mathcal{A}$ 0.515 0.806 0.540 ($\uparrow 0.025$) 0.824 ($\uparrow 0.018$)
UR-FUNNY $\mathcal{V},\mathcal{A}$ 0.219 0.563 0.258 ($\uparrow 0.039$) 0.598 ($\uparrow 0.035$)
UR-FUNNY $\mathcal{V},\mathcal{T}$ 0.037 0.204 0.038 ($\uparrow 0.001$) 0.202 ($\downarrow 0.002$)
DreamBooth + AVE + AudioSet $\mathcal{V},\mathcal{T},\mathcal{A}$ 0.514 0.804 0.533 ($\uparrow 0.019$) 0.816 ($\uparrow 0.012$)
@@ 5 Conclusions
In this paper, we analyze the limitations of fixed-anchor-binding (FABind) methods, including their over-reliance on a single anchor modality and their inability to capture both intra-modal and shared information among non-anchored modalities. To address these shortcomings, we propose adaptive anchor binding methods such as CentroBind, which align multi-modal embeddings to centroid-based anchors, removing the need for a fixed anchor modality. Our approach generalizes and extends methods like ImageBind. We also provide a theoretical analysis showing that CentroBind effectively captures both intra-modal and shared inter-modal information. Experiments on synthetic and real-world datasets demonstrate that CentroBind outperforms FABind across nearly all settings—including modality imbalance, varying backbone quality, differing numbers of modalities (including bimodal cases), and a wide range of dataset sizes—yielding a robust, unified representation space and validating our theoretical insights.
@@ A.1 Multi-modal learning
Typically, these models integrate uni-modal features extracted by modality-specific encoders (Seichter et al., 2021; Nagrani et al., 2021; Wu et al., 2022; Wang et al., 2020; Peng et al., 2022). For instance, (Madaan et al., 2024) introduces inter- and intra-modality modeling frameworks that treat the target as a composition of multiple modalities. Similarly, (Du et al., 2023) proposes a late-fusion approach for supervised multi-modal tasks, demonstrating that insufficient feature extraction from individual modalities negatively affects the model’s generalization ability. Additionally, (Zhang et al., 2024) addresses joint optimization by alternating between uni-modal learning scenarios and integrating modality-specific encoders with a unified head shared across all modalities.
Multi-view learning is closely related to multi-modal learning. Early work in (Tian et al., 2020a) studies multi-view representation learning through predictive and contrastive approaches across multiple modalities, demonstrating the efficacy of contrastive learning in multi-view settings. Building on these insights, numerous subsequent studies have integrated contrastive learning into multi-view tasks. For example, (Xing et al., 2019) proposes an approach that adaptively combines image-text representations for few-shot learning; however, this method still relies on labeled data. Some research leverages anchors or a fused modality as the anchor representation. (Zeng et al., 2021) introduces a unified prototype representation to address cross-modal retrieval imbalance by employing a reconstruction loss, but its reliance on unified prototypes as anchors restricts the capture of inter-view information across modalities. Meanwhile, (Huang et al., 2023) develops a tri-modal alignment pre-training task—extending text-video alignment to include a fused modality using pairwise contrastive learning; however, it does not explicitly handle intra-view learning. Furthermore, (Wang et al., 2022b) presents a cross-modal data augmentation technique for image-text multi-view learning, randomly replacing visually grounded words with diverse image patches to increase data variety and encourage token-level interaction across modalities. (Dufumier et al., 2024) introduces a contrastive multimodal approach that maximizes mutual information between augmented multi-modal features by effectively capturing redundant, unique, and synergistic interactions across modalities beyond traditional multi-modal alignment constraints. (Lin et al., ) presents inter-modality alignment combined with boundary expansion for multi-view classification to mitigate information redundancy. Nevertheless, this approach still overlooks intra-view information, indicating the need for methods that jointly consider both inter- and intra-view representations.
@@ A.2 Multi-modal alignment
Multi-modal learning addresses four key challenges (Liang et al., 2024c; Baltrušaitis et al., 2018; Liang et al., 2024d): managing interactions among redundant, unique, and synergistic features (Dumas et al., 2017; Liang et al., 2024a; b), aligning fine-grained and coarse-grained information (Wang et al., 2023; 2024a), reasoning across diverse features (Yang et al., 2023), and integrating external knowledge (Shen et al., 2022; Lyu et al., 2024). Among these challenges, multi-modal alignment is one of the core challenges that many researchers aim to solve.
A common method in multi-modal alignment is using cross-modal alignment by using attention mechanisms between pairwise modalities, such as vision-language (Tan & Bansal, 2019) and vision-language-audio (Tsai et al., 2019). Another effective approach is leveraging graph neural networks to align multi-modal datasets (Yang et al., 2021; Wilf et al., 2023). For instance, (Yang et al., 2021) transforms unaligned multi-modal sequence data into nodes, with edges capturing interactions across modalities over time. (Wilf et al., 2023) builds graph structures for each modality—visual, textual, and acoustic—and create edges to represent their interactions.
To enhance the generalizability of cross-modal representations, (Xia et al., 2024) employs a unified codebook approach, facilitating a joint embedding space for visual and audio modalities. Another prominent method achieves cross-modal alignment by leveraging large collections of image-text pairs, making it a widely adopted strategy in multi-modal learning (Radford et al., 2021; Zhang et al., 2022; Guzhov et al., 2022; Zhou et al., 2023).
A.3 Multi-modal domain generalization
Multimodal domain generalization (MMDG) aims to train models on data from multiple modalities (e.g., image, audio, text) and source domains, such that the models generalize to unseen target domains that share the same modalities. A primary challenge in MMDG lies in aligning heterogeneous modality-specific representations into a shared embedding space while preserving both shared semantics and modality-specific information.
(Fan et al., 2024) reduces generalization error by flattening the representation-space loss landscape using multi-modal feature interpolation and teacher-student distillation. This strategy mitigates modality dominance, in which stronger modalities (e.g., images) overpower weaker ones (e.g., audio), but it aligns representations only between modality pairs and assumes that all modalities are present during both training and inference. (Dong et al., 2023) decomposes modality features into shared and specific components and apply supervised contrastive learning to the shared part using label supervision. They further use cross-modal translation modules to reconstruct one modality’s representation from another. However, their framework depends on modality-specific translation paths and requires prior knowledge of available modalities, limiting its scalability to new combinations. (Dong et al., 2024) extends the setting to open-set domain generalization by introducing self-supervised tasks, masked cross-modal translation and multi-modal jigsaw puzzles, that enhance the model’s ability to detect unknown classes. While effective for class-level generalization, their method assumes a fixed modality set and does not support unseen modalities at test time.
In contrast, CentroBind constructs a unified and modality-agnostic embedding space by dynamically computing centroid-based anchors from all available modality embeddings within a sample. Instead of aligning only modality pairs or relying on fixed translation modules, CentroBind pulls each modality’s representation toward the centroid, enabling joint alignment without architectural changes. This approach supports any combination of available or unseen modalities and eliminates the need for predefined modality decomposition, making CentroBind scalable and adaptive for real-world multimodal scenarios.
A.4 Binding methods
@@ Appendix B Proofs
$\overset{\rm(b)}{=}\frac{M}{\tau}{\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,k})-|\mathcal{I}_{B}|\log\sum_{j\in\mathcal{I}_{B}}\exp\left(\frac{M{\boldsymbol{a}}_{k}^{\top}f_{i}({\boldsymbol{x}}_{i,j})}{\tau|\mathcal{I}_{B}|}\right)-|\mathcal{I}_{B}|\log C_{\mathcal{F},k,i}$
@@
[…]
@@
Appendix C Experiments
Algorithm 1 CentroBind
Initialize encoders $f_{1}^{(0)},f_{2}^{(0)},\cdots,f_{M}^{(0)}$.
for $t=0,1,\ldots,t_{\max}$ do
  Sample $B$ from multi-modal datasets $\{\mathcal{D}_{i}\}_{i}$.
  for $i=1,\ldots,M$ do
   Minimize $\mathcal{L}_{\rm CB}(f_{i}^{(t+1)}|\tau)$ in (8)
  end for
end for
In this section, we provide experimental details and additional results. Algorithmic expression of CentroBind is given in Algorithm 1.
@@ C.1 Experiments with synthetic datasets
$p_{\bf Z}({\boldsymbol{z}})$ $=\prod_{y=1}^{K}\pi_{y}\mathcal{N}({\boldsymbol{z}};\hbox{$\mu$}_{y},\hbox{$\Sigma$}_{y}),$ (26)
where $K$ is the number of mixture components, $\pi_{y}=\Pr({\bf Y}=y)$ is the component prior probability, and $\mathcal{N}({\boldsymbol{z}};\hbox{$\mu$}_{y},\hbox{$\Sigma$}_{y})$ denotes Gaussian PDF with mean $\hbox{$\mu$}_{y}\in\mathbb{R}^{d_{z}}$ and covariance matrix $\hbox{$\Sigma$}_{y}\in\mathbb{R}^{d_{z}\times d_{z}}$. This leads to the conditional PDF of $\bf Z$} as $p_{{\bf Z}|{\bf Y}}({\boldsymbol{z}}|y)=\mathcal{N}({\boldsymbol{z}};\hbox{$\mu$}_{y},\hbox{$\Sigma$}_{y})$.
@@
[…]
@@
The plot also reveals a crossover point where the loss curves intersect. We believe this occurs due to the number of InfoNCE losses optimized by CentroBind and FABind. Specifically, with $M$ modalities, CentroBind minimizes $M$ InfoNCE losses, while FABind minimizes M1M-1 InfoNCE losses. This results in a smaller loss for FABind when the encoders are well-trained, which explains the crossover point observed in Figure 6.
Temperature sensitivity.
We examine the effect of temperature parameter $\tau$ in InfoNCE loss by evaluating classification accuracy on the GMM synthetic dataset with different $\tau\in\{0.07,0.3,0.7\}$. Figure 7 displays the classification accuracy for each temperature setting. Three different $\tau$ yield similar performance gap between CentroBind and FABind , implying that the advantage of CentroBind is robust to hyperparameter. This further strengthens our claim that CentroBind overcomes the inherent limitation of FABind.
Complexity.
CentroBind introduces only one additional InfoNCE term per batch—the interaction between the learned centroid and the modality embeddings—so computational complexity rises merely from $O(M)$ to $O(M+1)$ for $M$ modalities. In memory, it stores a single extra vector (the current-batch centroid), adding a negligible constant overhead. Consequently, training and inference costs remain practically unchanged from FABind.
(a) $\tau=0.07$
(b) $\tau=0.3$
(c) $\tau=0.7$
Figure 7: Classification accuracy on GMM synthetic dataset. The embeddings are learned with different temperature parameter $\tau\in\{0.07,0.3,0.7\}$.
@@
Method Modality Sar-1 Spk-1 Spk-3 Spk-5 FABind $\mathcal{T}$ 0.606 0.219 0.458 0.632 UniBind 0.600 0.214 0.412 0.569 AudioCLIP 0.488 0.155 0.280 0.388 ViT-Lens 0.543 0.172 0.342 0.472 CentroBind 0.667 0.287 0.507 0.642 FABind $\mathcal{V}$ 0.668 0.375 0.587 0.691 UniBind 0.658 0.381 0.641 0.770 AudioCLIP 0.504 0.110 0.275 0.414 ViT-Lens 0.697 0.586 0.738 0.797 CentroBind 0.670 0.380 0.609 0.726 FABind $\mathcal{A}$ 0.639 0.201 0.457 0.599 UniBind 0.633 0.272 0.528 0.691 AudioCLIP 0.525 0.158 0.343 0.454 ViT-Lens 0.686 0.396 0.664 0.8 CentroBind 0.616 0.234 0.461 0.609 FactorCL $\begin{subarray}{c}\mathcal{V},\mathcal{A},\mathcal{T}\\ (\mathcal{V},\mathcal{A},\mathcal{T})\end{subarray}$ 0.699 - - - SimMMDG 0.725 - - - FABind 0.678 0.343 0.554 0.677 UniBind 0.646 0.383 0.622 0.764 AudioCLIP 0.530 0.119 0.261 0.378 ViT-Lens 0.731 0.506 0.736 0.812 CentroBind 0.704 0.346 0.594 0.733
@@
[…]
@@
[…]
@@
[…]
C.2 Experiments on MUStARD